# Shuchang Zhang, Hui Zhang, Hongxia Wang

### v2

@@ 3.2 Convergence rate
In this section, we will give the result of the sub-linear convergence rate of the gap between two successive iterations. We now present a technical Lemma [36] for convergence rate.
@@
Let $M>0$. Assume that $\{a_{k}\}_{k\in\mathbb{N}}$ is a sequence of nonnegative real numbers which satisfy $a_{1}<M$ and
@@
where $\gamma\in(0,1]$, $\{b_{k}\}_{k\in\mathbb{N}}$ is a sequence which is defined by $b_{k}=\min\{\frac{2}{\gamma k},1\}$, $\{a_{k}\}_{k\in\mathbb{N}}$ is a sequence of real numbers such that $c_{k}\leq M<\infty$. Then, the sequence $\{a_{k}\}_{k\in\mathbb{N}}$ satisfies
$a_{k}\leq\frac{MJ}{\gamma k},k\geq 1,$
@@
Corollary 3.7.
Let $\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$ be a $\mathcal{M}$-FNE operator, $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz, and $\{u^{k}\}$ be the sequence generated by (5). If $\mu_{k}=\min\{\frac{2}{k},1\}$, then
@@
    $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq\frac{2M^{\prime}}{k}$, i.e., $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}=\mathcal{O}(\frac{1}{k})$;
@@
Proof 3.8.
For (i), denote $M^{\prime}=\sup_{k\in\mathbb{N}}\{\left\|\mathbf{u}^{0}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}\}$, since
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+\left|\mu_{k+1}-\mu_{k}\right|M^{\prime},$
it derives (i) by Lemma 3.6. For (ii), since
@@
$\leq\frac{2M^{\prime}C}{k-1}+\frac{2M^{\prime}}{k(k+1)},k\geq 2.$
and thus $\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|=\mathcal{O}(\frac{1}{k})$.
3.3 HPPP with TV regularization
Chambolle-Pock(CP) is a first-order primal-dual algorithm to solve non-smooth convex optimization problems with known saddle-point structure [10]. Take the starting point $(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$, the CP primal-dual algorithm is written into the following form:
(14) $\left\{\begin{array}[]{ll}\mathbf{x}_{k+1}&=(I+\tau\partial f)^{-1}(\mathbf{x}_{k}-\tau\mathbf{K}^{*}\mathbf{y}_{k}),\\ \mathbf{y}_{k+1}&=(I+s\partial g^{*})^{-1}\left(\mathbf{y}_{k}+s\mathbf{K}(2\mathbf{x}_{k+1}-\mathbf{x}_{k})\right),\end{array}\right.$
here we consider TV regularization $\mathbf{K}=\nabla$. The convergence of CP algorithm was proved when step sizes satisfy $\tau s\left\|\mathbf{K}\right\|^{2}<1$. By introducing the operator $\mathcal{A}$ and preconditioner $\mathcal{M}$, set
(15) $\mathcal{A}=\begin{pmatrix}\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&\partial g^{*}\end{pmatrix},\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-\mathbf{K}^{*}\\ -\mathbf{K}&\frac{1}{s}I\end{pmatrix}.$
As fistly noticed in [20, 5], the CP method is a special PPP method $0\in\mathcal{A}\mathbf{x}^{k+1}+\mathcal{M}(\mathbf{x}^{k+1}-\mathbf{x}^{k})$. Bredies et al. proved the weak convergence of the CP algorithm about the degenerate case $\tau s\left\|\mathbf{K}\right\|^{2}=1$ [5], i.e., $\mathcal{M}$ is a semi-definite preconditioner. To apply HPPP with TV regularization, under the degenerate case $\tau s\left\|\mathbf{K}\right\|^{2}=1$ we obtain
(16) $\left\{\begin{array}[]{ll}\mathbf{x}_{k+1}&=\mu_{k+1}\mathbf{x}_{a}+(1-\mu_{k+1})(I+\tau\partial f)^{-1}(\mathbf{x}_{k}-\tau\mathbf{K}^{*}\mathbf{y}_{k}),\\ \mathbf{y}_{k+1}&=\mu_{k+1}\mathbf{y}_{a}+(1-\mu_{k+1})(I+s\partial g^{*})^{-1}\left(\frac{2s\mathbf{K}(\mathbf{x}_{k+1}-\mu_{k+1}\mathbf{x}_{a})}{1-\mu_{k+1}}-s\mathbf{K}\mathbf{x}_{k}+\mathbf{y}_{k}\right),\end{array}\right.$
where $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a}),(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$ are the anchor and initial points, respectively, and $\mu_{k}$ satisfies conditions(i)-(iii) of Theorem 3.1.
3.4 HPPP with implicit regularization
Recently, RED frameworks have achieved state-of-the-art performance for inverse problems by utilizing CNN (convolutional neural network) denoiser as regularization [14], such as DnCNN [46]. Ryu et al. proposed a nonexpansive residual of DnCNN to ensure the convergence of PnP methods [35]. Gradient step denoisers, defined as $D_{\sigma}=I-\nabla g_{\sigma}$ for Gaussian noise level $\sigma$, have been proposed to generalize RED [13, 24], where $g_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}$ is a scalar function parameterized by a differentiable neural network. The gradient step denoiser can be interpreted as the proximal operator of an implicit regularization [25].
@@ 3.4 HPPP with implicit regularization
The primal-dual algorithm to solve (4) is viewed as the fixed-point iteration $\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}\mathbf{u}^{k}$ with $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&(\partial g)^{-1}\end{pmatrix},\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-\mathbf{K}^{*}\\ -\mathbf{K}&\frac{1}{s}I\end{pmatrix}$ to solve (3). According to subsection 2.2, $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, we consider the special application $g=\phi,\mathbf{K}=I$ with the degenerate $\mathcal{M}(\tau s=1)$.
The proposed HPPP algorithm has the advantage of converging to a particular solution than PPP. Due to the data adaptability of CNN, which has the powerful expressiveness to represent the implicit regularization $\phi$ [44, 45]. Based on HPPP and GraRED, we propose the implicit Gradient RED via HPPP called GraRED-HP3 (see Algorithm 1). According to (8), we further discuss GraRED-P3 Algorithm 2 for IR problems.
PnP-ADMM [39, 11, 35] is a well-known optimization method to solve (1) when $g=\phi$, i.e.,
(17) $\mathbf{z}^{k+1}=$ $\mathrm{prox}_{\phi}(\mathbf{v}^{k}-\bar{\mathbf{u}}^{k})=D_{\sigma}(\mathbf{v}^{k}-\bar{\mathbf{u}}^{k}),$
$\mathbf{v}^{k+1}=$ $\mathrm{prox}_{\lambda f}(\mathbf{z}^{k+1}+\bar{\mathbf{u}}^{k}),$
$\bar{\mathbf{u}}^{k+1}=$ $\bar{\mathbf{u}}^{k}+\mathbf{z}^{k+1}-\mathbf{v}^{k+1}.$
Let $\mathbf{w}^{k}=\mathbf{v}^{k}+\bar{\mathbf{u}}^{k}$, then PnP-ADMM can be written into equivalent Douglas-Rachford Splitting (DRS) form:
(18) $\mathbf{w}^{k+1}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k}).$
Using $\mathbf{w}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k},\tau=s=1,\lambda_{k}=1$ in GraRED-P3 results exactly in the above DRS iteration. Thus, PnP-ADMM can be obtained from the perspective of PPP [6, 5], and its convergence follows by [5, Corollary 2.15].
The following Theorem 3.9 verifies the convergence of Algorithm 2 and Algorithm 2.
Theorem 3.9.
Assume that a denoiser $D_{\sigma}:\mathcal{H}\to\mathcal{H}$ satisfies homogeneity and Assumption (A), and has symmetric Jacobian, relaxing parameter $20<\inf_{k}\lambda_{k}\leq\sup_{k}\lambda_{k}<$ 2 and RED is convex, then $\{\mathbf{x}^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 2 converges to some uncertain solution $\mathbf{x}^{*}$ of (1), while $\{\mathbf{x}^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 1 converges to the particular solution $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$ of (3).
Proof 3.10.
Followed by [5, Corollary 2.10] and Corollary 3.3.
@@
1:initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, anchor point $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, diminishing anchor coefficient $\mu_{k}\in[0,1],\mu_{k}\to 0(k\to\infty),\sum_{k=0}^{\infty}\mu_{k}=+\infty$, and $R=I-D_{\sigma}$ is a residual of denoiser $D_{\sigma}$, degenerate preconditioner $\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-I\\ -I&\frac{1}{s}I\end{pmatrix},\tau s\leq 1$, where $I$ is identity matrix.
@@
4:     $\mathbf{x}^{k+1}=\mu_{k}\mathbf{x}_{a}+(1-\mu_{k})\mathbf{d}^{k}$
5:     $\mathbf{v}^{k+1}=R\left(s(2\mathbf{d}^{k}-\mathbf{x}^{k})+\mathbf{y}^{k}\right)$
6:     $\mathbf{y}^{k+1}=\mu_{k+1}\mathbf{y}_{a}+(1-\mu_{k+1})\mathbf{v}^{k+1}$
7:end for
8:$\mathbf{x}^{N}$.
@@
1:initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, relaxing parameter $\lambda_{k}\in[0,2]$, $R=I-D_{\sigma}$ is a residual of denoiser $D_{\sigma}$, $\mathcal{A}=\begin{pmatrix}\lambda\partial f&{I}\\ -I&\partial{\phi}^{*}\end{pmatrix}$, degenerate preconditioner $\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-I\\ -I&\frac{1}{s}I\end{pmatrix},\tau s\leq 1$, where $I$ is identity matrix.

### v4

@@ 3.2 Convergence rate
Based on the following technical Lemma, Sabach and Shtern [50] first showed that, for a general form of Halpern iteration, the fixed-point residual converges at a rate of $\mathcal{O}(1/k)$.
@@
Let $M_{1}>0$. Assume that $\{a_{k}\}_{k\in\mathbb{N}}$ is a sequence of nonnegative real numbers such that $a_{1}<M_{1}$ and
@@
where $\gamma\in(0,1]$, the sequence $\{b_{k}\}_{k\in\mathbb{N}}$ is defined by $b_{k}=\min\{\frac{2}{\gamma k},1\}$, and $\{c_{k}\}_{k\in\mathbb{N}}$ is a sequence of real numbers such that $c_{k}\leq M_{1}<\infty$. Then the sequence $\{a_{k}\}_{k\in\mathbb{N}}$ satisfies
$a_{k}\leq\frac{M_{1}J}{\gamma k},k\geq 1,$
@@
By applying the Lemma 3.9, we will establish a sublinear convergence rate for $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|$ and fixed point residual $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|$ or $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$.
Proposition 3.10.
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{u^{k}\}$ be the sequence generated by (5). If $\mu_{k}=\min\{\frac{2}{k},1\}$, then
@@
    $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$, i.e., $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}=\mathcal{O}(\frac{1}{k})$.
@@
Proof 3.11.
(i)According to (30), we have
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M\left|\mu_{k+1}-\mu_{k}\right|,$
where $M=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1},\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$, and $C_{1}$ is defined by (27). Applying Lemma 3.9 together with $c_{k}=M_{1}=M,J=2,\gamma=1,a_{k}=\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}$ and $\mu_{k}=b_{k}$ to the above inequality, we have $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$.
(ii)By (33), when $k\geq 2$ we have
@@
$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}=\mathcal{O}(\frac{1}{k}),$
where the second inequality follows from (i) and the fact that $\mu_{k}-\mu_{k+1}=\frac{2}{k(k+1)}(k\geq 2)$, and $M^{\prime}=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+CC_{1}$, and $C,C_{1}$ are defined in (25) and (27), respectively.
As for the second result, by the triangle inequality and $\mathbf{u}^{k+1}-\mathcal{T}\mathbf{u}^{k}=\mu_{k+1}(\mathbf{a}-\mathcal{T}\mathbf{u}^{k})$, for $k\geq 2$,
$\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|$ $\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\left\|\mathbf{u}^{k+1}-\mathcal{T}\mathbf{u}^{k}\right\|$
$\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|$
$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}+\frac{2M^{\prime}}{k+1}=\mathcal{O}(\frac{1}{k}).$
Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ according to [9, Proposition 2.3]. By [9, Theorem 2.13], the operator $\tilde{\mathcal{T}}=\mathcal{C}^{*}(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}$ is FNE, let $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}$ and $\mathbf{a}=\mathbf{u}^{0}$, then HPPP is equivalent to the following reduced algorithm:
(16) $\mathbf{w}^{0}=\mathcal{C}^{*}\mathbf{u}^{0},\mathbf{w}^{k+1}=\mu_{k+1}\mathbf{w}^{0}+(1-\mu_{k+1})\tilde{\mathcal{T}}\mathbf{w}^{k}.$
By [36, Theorem 2.1] and [52, Proposition 2.9], we further give the following (tight) optimal convergence rate for HPPP.
Proposition 3.12.
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$ be the sequences generated by (5) and (16). If $\mu_{k}=\frac{1}{k+1}$ and $\mathbf{a}=\mathbf{u}^{0}$, then
  1. (i)
    $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|\leq\frac{2\left\|\mathbf{w}^{0}-\mathbf{w}^{*}\right\|}{k+1}$ for $k\geq 0$ and $\mathbf{w}^{*}\in\mathrm{Fix}(\tilde{\mathcal{T}})$.
  2. (ii)
    $\left\|\mathbf{u}^{k}-{\mathcal{T}}\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq\frac{2\left\|\mathbf{u}^{0}-\mathbf{u}^{*}\right\|_{\mathcal{M}}}{k+1}$ for $k\geq 0$ and $\mathbf{u}^{*}\in\mathrm{Fix}({\mathcal{T}})$.
3.3 GraRED-HP3
@@
The primal-dual algorithm to solve (4) is viewed as the fixed-point iteration $\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}\mathbf{u}^{k}$ with $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&\partial g^{*}\end{pmatrix},\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-\mathbf{K}^{*}\\ -\mathbf{K}&\frac{1}{s}I\end{pmatrix}$. Under the degenerate case $\tau s\left\|\mathbf{K}\right\|^{2}=1$, the HPPP iteration is given by
(17) $\left\{\begin{array}[]{ll}\mathbf{d}^{k}&=(I+\tau\lambda\partial f)^{-1}(\mathbf{x}^{k}-\tau\mathbf{K}^{*}\mathbf{y}^{k}),\\ \mathbf{x}^{k+1}&=\mu_{k+1}\mathbf{x}_{a}+(1-\mu_{k+1})\mathbf{d}^{k},\\ \mathbf{y}^{k+1}&=\mu_{k+1}\mathbf{y}_{a}+(1-\mu_{k+1})(I+s\partial g^{*})^{-1}\left(\mathbf{y}^{k}+s\mathbf{K}(2\mathbf{d}^{k}-\mathbf{x}^{k})\right),\end{array}\right.$
where $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a}),(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$ are the anchor and initial points. Under the ideal conditions of RED and with Assumption (A), according to Lemma 2.6, both the denoiser $D_{\sigma}$ and the residual $R$ can be expressed as the proximal operators of implicit regularizers, i.e., $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}),R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$. Let $g(\mathbf{x})=\frac{\phi(s\mathbf{x})}{s}$, then $\phi^{*}=sg^{*}$ [5, Theorem 4.14], we can replace $(I+s\partial g^{*})^{-1}$ with the residual $R$ in (17). Set $\mathbf{K}=I$, the implicit Gradient RED via HPPP called GraRED-HP3 is proposed in Algorithm 3. Even if the conditions of RED are difficult to meet [45], recent developments [49, 54, 42, 21, 15, 10, 57] can learn FNE denoisers. According to Lemma 2.8, both the denoiser and the residual still act as the proximal operators of implicit regularizers.
3.3.1 Acceleration for PnP methods
PnP-ADMM (see Algorithm 1) is a well-known method to solve (1), which can be written into an equivalent DRS form [49, 33], i.e., $\mathbf{w}^{k+1}=\tilde{\mathcal{T}}\mathbf{w}^{k}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})$. When $\mathbf{w}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k},\tau=s=1,\mu_{k}=0$ in GraRED-HP3 results exactly in the DRS iteration. Thus, PnP-DRS is a special case of GraRED-HP3, which can be obtained from the perspective of PPP [11, 9]. The related GraRED-P3 algorithm is shown in Algorithm 4. Furthermore, Proposition 3.12 demonstrates that GraRED-HP3 is an accelerated PnP method achieving an $\mathcal{O}(1/k)$ convergence rate for the fixed-point residual.
Remark 3.13.
In [28], a worst-case convergence rate of $\mathcal{O}(\frac{1}{\sqrt{k}})$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$ was established for DRS, which also applies to PnP-ADMM. In contrast, GraRED-HP3 obtain an accelerated convergence rate of $\mathcal{O}(\frac{1}{k})$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$.
@@ 3.3.1 Acceleration for PnP methods
1:Initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, anchor point $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, and $R=I-D_{\sigma}$ is the residual.
@@
4:     $\mathbf{x}^{k+1}=\mu_{k+1}\mathbf{x}_{a}+(1-\mu_{k+1})\mathbf{d}^{k}$
5:     $\mathbf{y}^{k+1}=\mu_{k+1}\mathbf{y}_{a}+(1-\mu_{k+1})R\left(\mathbf{y}^{k}+s(2\mathbf{d}^{k}-\mathbf{x}^{k})\right)$
6:end for
7:$\mathbf{x}^{N}$.
@@
1:Initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, relaxing parameter $\lambda_{k}\in[0,2]$, and $R=I-D_{\sigma}$ is the residual.
@@
Algorithm 5 Relaxed RED-PRO via SD
1:Input: $\mathbf{x}^{0}\in\mathbb{R}^{n}$, $\{t_{k}\}_{k\in\mathbb{N}}$, $\alpha\in(0,1)$, $\lambda$, $\mu$, $J$, $N,\delta>0$ and the denoiser $T$.
2:for $k=0,1,2,\ldots,N-1$ do
3:     $\mathbf{x}^{k,0}=\mathbf{x}^{k}$
4:     for $j=0,1,2,\ldots,J-1$ do
5:         $\mathbf{x}^{k,j+1}=T_{\alpha}\left(t_{j}\mathbf{x}^{k}+(1-t_{j})\mathbf{x}^{k,j}\right)$
6:     end for
7:     $\mathbf{v}^{k}=\frac{\delta}{\|\mathbf{x}^{k}-\mathbf{x}^{k,J}\|}\mathbf{x}^{k}+\left(1-\frac{\delta}{\|\mathbf{x}^{k}-\mathbf{x}^{k,J}\|}\right)\mathbf{x}^{k,J}$
8:     $\mathbf{x}^{k+1}=\mathbf{x}^{k}-\mu\left(\nabla f(\mathbf{x}^{k})+\lambda(\mathbf{x}^{k}-\mathbf{v}^{k})\right)$
9:end for
10:Output: $\mathbf{x}^{k+1}$
3.3.2 Compared with RED-PRO
Cohen et al. proposed the RED-PRO model [20],
(18) $\min_{\mathbf{x}\in\mathrm{Fix}(T)}\frac{1}{2\sigma^{2}}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|^{2}.$
RED-PRO sheds new insights on denoiser prior, which theoretically bridges between RED and PnP prior. However, practical denoisers often have narrow fixed-point sets, leading to suboptimal recovery solutions. To address this issue, they relax the hard constraint of $\mathrm{Fix}(T)$ and replace it with a dilated fixed-point set, defined for some $\delta>0$ as
$B_{\delta}(T)=\left\{\mathbf{x}\in\mathbb{R}^{n}:\left\|\mathbf{x}-P_{\mathrm{Fix}(T)}(\mathbf{x})\right\|\leq\delta\right\},$
where $P_{\mathrm{Fix}(T)}(\mathbf{x})$ is the fixed-point projection of $\mathbf{x}$ onto $\mathrm{Fix}(T)$. The relaxed RED-PRO (RRP) problem is formulated as
(19) $\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{1}{2\sigma^{2}}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|^{2}+\frac{\lambda}{2}\left\|\mathbf{x}-P_{B_{\delta}(T)}(\mathbf{x})\right\|^{2}.$
Algorithm 5 outlines RRP, where $T_{\alpha}=\alpha T+(1-\alpha)I$ with the demi-contractive denoiser $T$. When the steepest descent method solves (19), it requires an inner loop to calculate the fixed-point projection $P_{\mathrm{Fix}(T)}(\mathbf{x}^{k})$, and the line 5 in RRP is equivalent to the Halpern iteration.
Comparing Algorithm 3 with the inner loop of Algorithm 5, both methods share the same algorithmic form of Halpern iteration and utilize identical anchor coefficients. In both cases, finding $P_{\mathrm{Fix}(T)}(\mathbf{x}^{k})$ or $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$ corresponds to a bilevel optimization problem that finds the point in the fixed-point set closest to the anchor point $\mathbf{x}^{k}$ or $\mathbf{a}$.
The key difference is that HPPP extends Halpern iteration to a degenerate form by employing a positive semi-definite metric $\mathcal{M}$, whereas the denoiser operator $T$ in RED-PRO is defined under the standard metric $\mathcal{M}=I$. From an optimization perspective, the RRP algorithm is based on gradient descent and uses the classic Halpern iteration in its inner loop. In contrast, $\mathcal{T}$ in GraRED-HP3 is an algorithmic operator and has no inner loop. From a PPP standpoint, we integrate HPPP with denoiser priors to propose Algorithm 3.